2^{nd} EUROPEAN CONGRESS OF MATHEMATICS Budapest, Hungary July 21–27, 1996

Budapest welcomes you

SECOND ANNOUNCEMENT

Bures-sur-Yvette, December 4th 1995

Dear Colleagues and Friends,

It is my privilege to encourage you to participate in the second European Congress of Mathematics to be held next July, and to enjoy the deep sense of culture and warm hospitality of Hungary, the host country. This big event has been placed under the auspices of the European Mathematical Society since the Council of the society decided to continue the series inaugurated in Paris four years ago.

Allow me to express some of my feelings towards this endeavour. A scientific congress is in the first instance a great occasion for meeting people and for listening to distinguished colleagues. It also offers a special opportunity for assessing where one stands. In the context of such a European demonstration, I feel it is of the utmost importance to cultivate and confront our diversity. The Round Tables, one of the distinctive features of the European congresses, allow us to make our differences visible and to strengthen our efforts to prevent them from becoming obstacles. This will certainly make us more receptive to the theme under which the Scientific Committee has decided to place the Congress, the Unity of Mathematics.

The success of the Congress will above all be the result of the hard work of our Hungarian colleagues. It will to a lesser extent reflect our personal involvement in its preparation and its oganization. It is up to us to make the European Mathematical Society an efficient tool for developing projects of this kind, with the general public of mathematicians in mind, or an even wider one. Indeed, we cannot think of ourselves as alone in a corner of the planet. It is our task to make Mathematics accessible to the public at large, and to try and understand its power and limitations. A congress like this one provides an exceptional opportunity.

Europe can only prosper in developing compatibility between its inhabitants, no matter what their speciality. This is a condition for the passage of its existence from a virtual world of symbols to the real one. It falls to us to work at this, and we hope to succeed.

Looking forward to meeting you in Budapest in July, I remain

Sincerely yours,

Jean-Pierre Bourguignon President of the European Mathematical Society

The Organizing Committee extends a hearty invitation to all members of the international mathematical community to attend the Second European Congress of Mathematics in Budapest, Hungary, July 21–27, 1996.

The Congress is organized by the János Bolyai Mathematical Society and will be held under the auspices of the European Mathematical Society.

1. LOCATION

Budapest, the capital of Hungary is situated on the banks of Danube. The city is a major attraction in world tourism, offering rich programmes in culture, sport and nightlife as well as spectacular walking tours of the city of Pest and of the nearby hills of Buda.

Events will be at three convenient locations:

(i) the modern Convention Center providing 1700 seats for the plenary lectures, as well as closed circuit television for remote viewing,

(ii) the lecture rooms of the Technical University, where all other mathematical events will take place, (iii) the Hungarian National Gallery in the central building of the Royal Palace on the Castle Hill of

Buda, where the Banquet is to be held.

2. SCIENTIFIC PROGRAM

2 .1. Opening Ceremony, Monday, July 22, 10:00 am at the Convention Center, including

the European Mathematical Society prize awards and followed by a plenary lecture at 11:30.

2 .2. Plenary Lectures, 50 minutes long, are scheduled in the Convention Center on Monday, July 22 at 11:30 am, 2:30 pm, 3:30 pm, 4:30 pm; Wednesday, July 24 at 9:00 am, 10:00 am, 11:00 am; Friday, July 26 at 9:00 am, 10:00 am, 11:00 am. It is our pleasure to announce the following speakers:

N. Alon, G. Ben Arous, D. McDuff, B. Dubrovin, J. Kollár, J. Laskar, A. Merkurjev, V. Milman, S. Müller, J.-P. Serre .

2 .3. Parallel sessions of 45 minute talks are scheduled at the Technical University on Tuesday, July 23 from 9:00 am to 5:00 pm; Wednesday, July 24 from 2:00 pm to 5:00 pm; Thursday, July 25 from 9:00 am to 5:00 pm; Friday, July 26 from 2:00 pm to 4:00 am. The list of speakers will be made public on the EMS server whose address is http://www.emis.de and by anonymous ftp from circle.math-inst.hu:/pub/ecm2.

2 .4. Round tables covering 7 different topics are scheduled at the Technical University on Tuesday, July 23, 9 am – 12 am, and 2 pm – 5 pm; Thursday, July 25, 9 am – 12 am, and 2 pm – 5pm. Round table topics (and chairs if known already): A) Communication in mathematics - J. Körner; B) Mathematical games - D. Singmaster; C) Demography of mathematicians; D) Women and mathematics - K. Haag; E) Public image of mathematics; F) Mathematics and Eastern Europe - D. Cioranescu; G) The role of publication in the career.

2 .5. Award Lectures. EMS prize winners will be given the opportunity of giving a 30 minute

presentation of their work.

2 .6. Poster sessions. All registered members may present their work in scheduled poster

sessions provided they submit an abstract together with the AMS classification code(s) to the János Bolyai Mathematical Society (Fő u. 68, H-1027 Budapest, Hungary) by March 15, 1996. Abstracts sent by e-mail will not be considered.

Authors are responsible for the contents of their poster and are requested to prepare and bring

their sheets ready for presentation. Posters must be of size A2 or smaller.

Poster sessions are scheduled at the Technical University, Tuesday, July 23 from 9:00 am to 5:00 pm; Wednesday, July 24 from 2:00 pm to 5:00 pm; Thursday, July 25 from 9:00 am to 5:00 pm; Friday, July 26 from 2:00 pm to 4:00 am.

2 .7. Informal discussion. Lecture rooms and additional premises at the Technical University

will be available from 9:00 am - 9:00 pm for free dicussions.

2 .8. Movies of mathematical interest, films and videos on mathematicians will be shown.

Participants may suggest specific films for presentation. Such requests must be filed to the Organizing Committee by March 15 and a copy of film be sent not later than May 15.

2 .9. Exhibitions (including books).

2 .10. Closing Ceremony will take place in the Convention Center on July 26 at 7:00 pm,

followed by the Amadinda concert.

2 .11. A listing of satellite conferences is enclosed on a separate leaflet.

3. PUBLICATIONS

All registered participants will receive a full program, the Abstracts of Lectures, and a list of participants (registered prior to July 1, 1996) upon arrival at the Congress Reception Desk.

The Proceedings of the Congress can be ordered at a special discount price during the meeting. The Proceedings will contain refereed papers of lectures and summarized discussions of Round Table

meetings.

4. IMPORTANT DEADLINES

4.1. Grant applications received after January 31, 1996, will not be processed. 4.2. Poster proposals and early registrations are to be received by March 15, 1996. 4.3. Hotel reservations are guaranteed upon requests received by June 1, 1996.

5. SOCIAL AND CULTURAL EVENTS

1. All registered participants are welcome to the Banquet of the Congress held on July 22 (Monday) at 7.30 p.m. in the premises of the National Gallery.

2. A concert of unique musical experience will be presented by the Amadinda Percussion Group on July 26 (Friday) at the Convention Center following the Closing Ceremony. The concert is free of charge for all registered participants.

3. Additional cultural programmes. Various concerts and folklore programmes will be offered every evening. The Third Anouncement will contain a complete list of these optional events. The list and prices will also be available at the Congress Reception Desk.

6. REGISTRATION to the Congress

Applicants for registration should complete and return the enclosed Registration Form. All participants, including the invited speakers, are to submit their completed Registration Form to the

Registration Office.

All applications should be accompanied by the full payment of the registration fee. Registration Office: MALÉV AIR TOURS, Roosevelt tér 2, H-1051, Budapest Hungary Tel: +361-266 7836, Fax: +361-266 6006, e-mail: ecm2mat@math-inst.hu.

Registration fee in Swiss Francs if received:

prior to March 15  March 16 – May 15  after May 16 and on site
Members of EMS  130  165  240
Non-members  165  200  240
Accompanying persons  110  130  160

For way of payment see below.

On-site registration and reception desks

On-site registration as well as the participants’ registration packages and additional information will be available at the reception desks in the Assembly Hall of the Technical University, building “K”, main entrance. Reception service will be provided from July 21, Sunday to July 27, Saturday at the following hours:

July 21, 14.00 – 22.00; July 22-26, 8.00 – 18.00; July 27 8.00 – 14.00.

The registration fee includes

– admission to all sessions of the congress;

– admission to the Banquet on July 22;

– admission to the exhibitions and films;

– admission to the Amadinda concert on July 26;

– ticket for free transport in Budapest;

– the registration package.

The registration fee of accompanying persons includes

– admission to the Banquet on July 22;

– admission to (i) a sightseeing tour or (ii) a guided museum tour (see the optional tours);

– admission to the exhibitions and films;

– admission to the Amadinda concert on July 26;

– ticket for free transport in Budapest.

7. PAYMENT

Payment of the registration fee, deposits on accommodation and optional tours can be made in any major currency (using the rate of exchange on the day of payment) by certified cheque or by any major credit card. No personal cheques will be accepted.

Certified cheques should be sent to: MALÉV AIR TOURS, Congress Department, Roosevelt tér 2, H – 1051 Budapest, Hungary

Please make sure that your name and program number 4006-06-21076 are clearly written on your cheque.

Cancellation of participation and refunds

Cancellation has to be made in writing to MALÉV AIR TOURS, Congress Department. The amount of the refund depends on the date of cancellation as follows.

cancellations before June 1: 90 % of the total amount,

cancellations between June 1 and July 1: 50 % of the total amount, cancellations after July 1: no refund.

Note: All refunds will be paid after the congress.

8. HOTEL INFORMATION

Malév Air Tours is ready to accept reservations of hotel rooms and will provide special reduced prices for participants.

Prices of rooms in some recommended hotels:

price in Swiss Franks of

category  name of hotel  single room  double room
*****  Hilton  245  275
Hyatt  170  205

Marriott  295  320
****  Novotel  150  170
Flamenco  136  162
Korona  162  180
Gellért  160  270
***  Taverna  126  165
**  Ventura  77  85
Platanus  77  85
Griff  47  60
Student hostel  20  24

Please complete the enclosed hotel reservation form and return it with your deposit by June 1, 1996 (see PAYMENT above). Receipt of the returned form will be confirmed and you will be notified of the hotel where the preliminary reservation has been made. Prices are quoted per night per room and include breakfast, service and all taxes. (Tax is subject to increase by government regulation.) Since the Congress takes place in the high tourist season, an early booking of hotel rooms is highly recommended.

Please note: the deadline for guaranteed room reservation is June 1, 1996.

Cancellation

Cancellation of hotel reservation can only be made by written notice sent to the MALÉV AIR TOURS Congress Department.

Refunds:

- - cancellations recieved prior to June 1, 1996, are subject to a 10% cancellation charge,

- - no refunds for cancellations received after June 1, 1996.

9. GRANTS for mathematicians from Eastern European countries

A very limited number of grants for such colleagues will be available. These grants should cover the registration fee and part of the student style lodging expenses.

Copies of the Second Announcement going to Eastern European countries have a special grant application form inserted. If you want to apply for such a grant you should carefully fill in this form and send it together with your registration form in time to reach ECM2 no later than January 31, 1996. If you apply for a grant, you need not pay the registration fee and the hotel deposit when you send in the application forms. Applicants for grants will be notified of the decision by April 25, 1996. From those applicants who do not get the support the prior-March-15 registration fee will be accepted if sent in till May 15.

Several requests for financial support have already reached the congress office during preregistration. These requests have not been processed. Therefore, even if you already have placed such a request, you should nonetheles follow the procedure outlined above. In effect, grants will be awarded on the basis of the documents to be enclosed with the grant request form.

10. OPTIONAL PROGRAMMES (tours and excursions)

All tours are based on a minimum number of 10 participants. If this number is not reached, alternate arrangements can be chosen, or the deposit will be fully refunded upon request at the MALÉV Air Tours Information Desk during the event. All tours are led by English speaking guides and depart from the main entrance of the Technical University.

( i) S/S – City Sightseeing Tour

This three-hour tour presents the most attractive features of Budapest.

Dates: July 21 and 22, Price: CHF 22.

( ii) BM – Budapest Museums

The tour will visit two of the most important museums: the National Museum and the National

Gallery in the Buda Castle.

Date: July 23, Price: CHF 25.

( iii) BN – Budapest by Night

Sightseeing in the evening with a sample of Budapest’s champagne night life.

Date: July 25, Price: CHF 35.

( iv) HD – Hungarian Dinner with Folklore Show

Dinner in a traditional Hungarian restaurant with a floorshow highlighting Hungarian folklore. Date: July 26, Price: CHF 50.

( v) DB – Excursion to the Danube Bend

One-day trip along the Danube visiting three historic river towns: Esztergom, Visegrád and

Szentendre. The tour includes a lunch.

Date: July 24, Price: CHF 60.

( vi) PU – Excursion to the Puszta

A taste of the traditional life of the horsemen of the Great Hungarian Plain. The horse show

will be followed by a lunch in a typical Hungarian inn.

Date: July 28, Price: CHF 68.

( vii) LB – Lake Balaton

Balaton, the “Hungarian Sea”, is the largest lake in Central Europe. The two-day tour visits

the nicest and most important places around the lake, and also the world famous Herend porcelain factory. Meals are included in the price.

Date: July 20-21,

price: CHF 315/person in double room, CHF 77/single supplement

Cancellation of optional tours

before June 1, 1996: total refund; between June 1 and July 1: 90 % of the payment; after July 1: no refund.

11. GENERAL INFORMATION

VISA Visitors from most European countries (and several others, including the USA) can enter Hungary

without a visa. In case of doubt please consult a Hungarian embassy or consulate. The congress office can issue an official invitation if needed.

CUSTOMS According to the present regulations visitors may not carry more than 1000 Hungarian Forints when

entering or leaving the country. There is no limitation on foreign currencies. 1 liter of hard liquor, 2 liters of wine and 2000 cigarettes can be brought into Hungary duty free.

CLIMATE The climate in Hungary is generally pleasant and sunny in July with an average temperature of 25

degrees Celsius (77 F).

CURRENCY EXCHANGE Exchange offices are available at the airport, at railway stations, in hotels, banks and during the confer-

ence at the registration desk. (Visitors are advised not to change money in the street.)

INSURANCE Health and luggage insurance in Hungary may not be included in your current policy. Insurance covering

the period of your visit to Hungary can be purchased from MALÉV AIR TOURS when you arrange your travel.

TRAVEL See the list of MALÉV representations on page ...

MEALS

There is a large variety of restaurants in any district of Budapest. Prices are low relative to Western European standards. Inexpensive lunch will be available at the restaurant of the Technical University.

COMMITTEES

The Scientific Committee consists of

J. Moser (CH), chair, L. Babai (H), E. Bayer (F), L. Carleson (S), C. De Concini (I), S.K. Donaldson (UK), B. Engquist (S), S. Hildebrandt (D), S. Novikov (Russia), E. Pardoux (F), A. Schinzel (PL), D. Szász (H).

The Round Table Committee consists of

B. Prum (F) chair, C. Duhamel (F), A. Figà-Talamanca (I), D. Gabay (F), A. Recski (H) co - chair H. Sendov (BG). The Satellite Meeting Committee is co-chaired by G.O.H. Katona and Á. Szendrei.

Please note that this announcement (as well as the later ones) is (will be) available by anonymous ftp from circle.math-inst.hu:/pub/ecm2 and at the Server of the European Mathematical Society

http://www.emis.de.

Important notice: If you have any questions about

- - – lodging and optional programmes, please contact MALÉV Air Tours, Congress Department, Roosevelt tér 2, H - 1051 Budapest, Hungary; fax: (36 - 1) 266 - 6006; e-mail ecm2mat@math-inst.hu;

- - – scientific program, exhibitions and grants, please contact J. Bolyai Mathematical Society, Fő u. 68, H - 1027 Budapest, Hungary; fax: (36 - 1) 201 - 6974; e-mail ecm2jbms@math-inst.hu.

Budapest, December 6, 1995.

G.O.H. Katona President of the Organizing Committee

A. Balog  Cecília Kulcsár  L. Márki
Secretary  Treasurer  Representative of the EMS

Enclosures